OpenStudy (superhelp101):

Baking soda (NaHCO3) is an active ingredient in some antacids. Determine the percent composition of oxygen in baking soda. 19% 21% 42% 57%

OpenStudy (superhelp101):

@Abhisar

OpenStudy (superhelp101):

@DW101796

OpenStudy (superhelp101):

@Hero

OpenStudy (superhelp101):

@sammixboo @Abhisar @dan815 @deshawn1 @dcknox123 @Caitlyn_Morgan @Luigi0210 @livelovechloeee @luiss_69

OpenStudy (anonymous):

42%

OpenStudy (abhisar):

Have u tried it ur self ?

OpenStudy (superhelp101):

no i have no clue how to do it

OpenStudy (abhisar):

First calculate the molecular mass of NaHCO\(_3\).

OpenStudy (abhisar):

Na is 22.99 H is 1.01 C is 12.01 O is 16.00

OpenStudy (abhisar):

so molecular mass will be ?

OpenStudy (superhelp101):

wait so am I to add all of these Na is 22.99 H is 1.01 C is 12.01 O is 16.00

OpenStudy (abhisar):

yes. Don't forget to add O three times. Because its O\(_3\)

OpenStudy (superhelp101):

52.01 what difference would it make to add 0 three times

OpenStudy (abhisar):

not 0, I said O (oxygen)

OpenStudy (abhisar):

Molecular mass will be 23+1+12+(16*3)=84

OpenStudy (abhisar):

got it ?

OpenStudy (abhisar):

I have rounded off the figures

OpenStudy (superhelp101):

yep. how do you know what the total mass of the element is

OpenStudy (abhisar):

Now, % composition = \(\huge\frac{Total~mass~of~element}{Molecular~mass~of~Compound}\) * 100

OpenStudy (abhisar):

=\(\huge\frac{16*3}{84}\) *100 = \(\huge\frac{48}{84}\) *100= 57%

OpenStudy (superhelp101):

I see but where is 48 coming from?

OpenStudy (abhisar):

Mass of one oxygen atom is 16 and there are 3 oxygen atoms in the compound. So total mass of oxygen element = 16*3=48

OpenStudy (superhelp101):

AHH I See Thanks for teaching mE!

OpenStudy (abhisar):

\(\color{red}{\huge\bigstar}\huge\text{You're Most Welcome! }\color{red}\bigstar\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)

OpenStudy (abhisar):

@dcknox123 Seeking/Giving direct/Just answers is against the \(\large\sf\color{red}{\underline{\href{ /code-of-conduct}{code~of~conduct}}}\) of openstudy.